Required information

[The following information applies to the questions displayed below.]

Adger Corporation is a service company that measures its output based on the number of customers served. The company provided the following fixed and variable cost estimates that it uses for budgeting purposes and the actual results for May as shown below:

Fixed Element
per Month
Variable Element per Customer Served Actual Total
for May
Revenue $ 5,600 $ 182,000
Employee salaries and wages $ 55,000 $ 1,600 $ 110,300
Travel expenses $ 850 $ 27,200
Other expenses $ 34,000 $ 32,600

When preparing its planning budget the company estimated that it would serve 30 customers per month; however, during May the company actually served 35 customers.

Required:

13. What amount of other expenses would be included in Adger’s planning budget for May?

14. What activity variance would Adger report in May with respect to its revenue? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

15. What activity variances would Adger report with respect to each of its expenses for May? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)
Answer #1
13
Other expenses 34000
14
Flexible budget Activity variance Planning budget
196000 28000 F 168000
15
Flexible budget Planning budget Activity variance
Employee salaries and wages 111000 103000 8000 U
Travel expenses 29750 25500 4250 U
Other expenses 34000 34000 0 None
Workings:
Revenue:
Flexible budget 196000 =35*5600
Planning budget 168000 =30*5600
Expenses: Flexible budget Planning budget
Employee salaries and wages =55000+(1600*35) =55000+(1600*30)
Travel expenses =850*35 =850*30